OpenStudy (anonymous):

Solve the equation. 6(t+11)+8(t-19)

OpenStudy (mathstudent55):

An equation needs an equal sign. Is this an equation with some of it missing, or is this simply an expression that needs to be simplified?

OpenStudy (anonymous):

oh yes sorry 6(t+11)=8(t-19)

OpenStudy (anonymous):

14 t-86 just expand the given like: 6 t + 11*6+ 8t-8*19 then add the varaibles alone then the constants

OpenStudy (mathstudent55):

Use the distributive property on both sides to get rid of parenthses: 6(t+11)=8(t-19) 6t + 66 = 8t - 152 Can you do the next step?

OpenStudy (anonymous):

Not sure how to do the next step

OpenStudy (mathstudent55):

Subtract 8t from both sides and subtract 66 from both sides. This will get all t's on left side and all numbers on right side.

OpenStudy (mathstudent55):

6t + 66 = 8t - 152 -8t - 66 -8t - 66 -----------------------

OpenStudy (mathstudent55):

6t - 8t = ? 66 - 66 = 0 8t - 8t = 0 -152 - 66 = ?

OpenStudy (anonymous):

is it 86?

OpenStudy (anonymous):

t=86

OpenStudy (mathstudent55):

6t + 66 = 8t - 152 -8t - 66 -8t - 66 ----------------------- -2t = -218 Divide both sides by -2: t = 109

OpenStudy (anonymous):

Hi Click the document :)

OpenStudy (anonymous):

oh ok i got it that makes sense thank you @mathstudent55 !

OpenStudy (anonymous):

Thank you @JUICEDECAFE

OpenStudy (mathstudent55):

@marixoxo76 What is the correct answer?

OpenStudy (anonymous):

No Problem :)